OpenStudy (anonymous):

A right triangle has an angle that measures 34, and the adjacent side measures 17. What is the length of the hypotenuse to the nearest tenth?

OpenStudy (anonymous):

I think the equasion should be 17cos34 because adjacent and hypotenuse is cos but I get 14.0936..... and that isn't one of my choices.

OpenStudy (campbell_st):

draw a picture  the question involves trigonometry

OpenStudy (anonymous):

My choices: 20.5 25.2 30.4 34

OpenStudy (campbell_st):

ok... slight problem with your solution. the hypotenuse is the longest side in the triangle. So the and your answer is shorter than the adjacent side.. so the ratio should be \[\cos(34) = \frac{17}{h}\] is that ok...?

OpenStudy (campbell_st):

all you do is swap the postions of cos(34) and h

OpenStudy (campbell_st):

then calculate \[h = \frac{17}{\cos(34)}\]

OpenStudy (anonymous):

Oh, okay. So it would be 20.5.

OpenStudy (anonymous):

Thank you. :)